OpenStudy (anonymous):

√3 × √7 a. √24 b. √63 c. 3√7 d. None of the above

OpenStudy (anonymous):

D option

OpenStudy (anonymous):

\[\sqrt{3} * \sqrt{7} = \sqrt{21}\]